# Towards Effective Multiple-in-One Image Restoration: A Sequential and Prompt Learning Strategy

### 07 Jan 2024 v1

@@ 1 Introduction
There are some “all-in-one”  [^1] We believe the term of “multiple-in-one” is more precise and appropriate than “all-in-one”. methods  [24, 38, 41] that attempt to handle multiple IR tasks with a single model. However, their setups and explorations have some limitations. Firstly, many of them only take 3 or 4 IR tasks into consideration. Such a small number of tasks cannot reflect the training conflict between different tasks. Secondly, many of the “all-in-one” methods simply adopt the datasets from individual IR tasks in training. However, the ground-truth (GT) images of different datasets have uneven image quality. For example, the GT images from Rain1200 [57] and RESIDE [23] datasets contain obvious JPEG artifacts. The low quality of GT images will mislead the training and evaluation of IR networks. These limitations affect the exploration of the research problems along this line. Therefore, despite the progress made by the above methods, by far it remains a challenging issue on how to train a single model to effectively handle multiple IR tasks.
To solve the above mentioned problem, we propose the formulation of Multiple-in-One (MiO) IR, which aims to tackle multiple IR tasks using a single model. There are two pivotal challenges in MiO IR: diverse objective optimization and task adaptation. We then develop two effective and complementary strategies – sequential learning and prompt learning – attempting to address these two challenges, respectively. Specifically, as illustrated in Fig. 1(a), we consider 7 popular IR tasks, including super-resolution, deblurring, denoising, deJPEG, deraining, dehazing and low-light enhancement, and train a single model to handle them all. Compared with the setting in previous works [24, 38, 41], the proposed MiO setup employs high-quality GT images to generate the training and testing data, avoiding the risk of low-quality supervision signals. This formulation also allows us to explore the unique challenges of MiO IR.
@@
For the challenge of task adaptation, we propose the prompt learning  [^2] Prompt learning is similar to conditional learning in this work. Please refer to the related work section for more discussions. strategy. An appropriate prompt can help the network better understand the task at hand and adjust the direction of reconstruction. As illustrated in Fig. 1(c), we provide two methods of prompt learning. One uses additional input as prompt to obtain the task type explicitly (like that in  [38, 41]), while the other adaptively extracts dynamic visual prompt from the input image. These two methods represent the two extreme cases of prompt learning, and are favourable to different application scenarios. As shown in Fig. 2, together with sequential learning, the explicit prompt learning improves the average PSNR by 0.84/1.21 dB for SRResNet/SwinIR, respectively, while the adaptive prompt learning achieves an improvement of 0.24/0.95 dB for SRResNet/SwinIR, respectively. It is worth mentioning that, unlike previous approaches [24, 59], our adaptive prompt learning strategy does not require any specially designed supervision, and its higher generalization ability can be witnessed by an average PSNR improvement of 1.07/0.62 dB for SRResNet/SwinIR across five out-of-distribution test sets. Besides, our strategies can also enhance the state-of-the-art method PromptIR [41] by 1.1 dB with only 75% of its parameters.
@@ 2 Related Work
Prompt learning is originally known from the research on how to introduce additional texts (i.e., prompts) as inputs to pre-trained large language models so that the desired outputs can be obtained [42, 3]. With further research, it becomes common to use different forms of prompts in model training or fine-tuning [2, 50]. In IR field, ProRes [38] and PromptGIP [31] employ additional input images or image pairs as prompts to tell the model what the IR task is. PromptIR [41] employs a classifier architecture to process additional degradation context. These methods can be viewed as explicit prompt learning. However, in real-world IR applications, sometimes it is difficult to explicitly assign an exact task type for the given image. So it is anticipated to extract information adaptively from the input image as prompt [24, 48]. In this work, both explicit prompt learning and adaptive prompt learning are investigated for the proposed MiO IR formulation.
@@ 3.3 Prompt Learning
For explicit prompt learning, we use additional input as prompt to tell the network what the task type is (like that in  [38, 41]). We fix the prompt parameters and bind them with different IR tasks, and then train a classifier to select the prompt for the corresponding IR task during training and testing. Since it is not diffucult to training a high-accuracy classifier for IR tasks, this prompt selection method can be viewed as informing the network the explicit task types directly (manual selection). Considering that the difficulty of explicit prompt learning is not very high, it is supposed to have good performance on in-distribution data.
For real-world low-quality images, sometimes even human subjects cannot explicitly tell the IR task type, and hence explicit prompt learning may fail in such cases. The adaptive prompt learning can be a remedy, where we use the input image as visual prompt to extract task type information adaptively. As illustrated in Fig. 1(c), it has the same architecture as explicit prompt learning except that the input image is used as prompt, instead of the additional input. Because we do not impose any additional constraints on the extractor of prompt, adaptive prompt learning models are much more difficult to train than explicit prompt learning. However, once learned, it has better out-of-distribution generalization capability because the network is able to decide what features to extract by itself.
@@ 4 Experiments and Analysis
In this section, we firstly present the implementation details of MiO IR, including training / testing data and training settings, in Sec. 4.1. Then we show the effectiveness of our proposed learning strategies on in/out of distribution test sets in Sec. 4.2, and use our proposed strategies to enhance the existing state-of-the-art method in Sec. 4.3. After that, we apply our strategies to more backbones in Sec.4.4, and interpret the effectiveness of our strategies from the perspective of degradation representation in Sec. 4.5. Finally, we summarize the results of common backbone networks on MiO IR in Sec. 4.6 and Sec. 4.7 for easy access and comparison with future methods.
@@ 4.2 Effectiveness of Learning Strategies
First, we evaluate the effectiveness of sequential learning on the In-Dis test group. As depicted in Tab. 3, compared with mixed learning (‘-M’), basic sequential learning (‘-S’) can enhance the average PSNR of SRResNet/SwinIR by 0.29/0.85 dB across the 7 tasks. It is worth mentioning that, ranging from 0.18 dB (SRResNet on denoising task) to 2.31 dB (SwinIR on deraining task), the performances on all the 7 tasks using both the two backbones are improved, even the least trained task (i.e., low-light enhancement).
@@
Table 5: PSNR results of state-of-the-art method PromptIR [41] with and without our learning strategy on In-Dis, Out-Dis and Uknown. All models are trained under our MiO IR formulation. Our sequential learning could directly enhance PromptIR. By replacing the original prompt method of PromptIR with our EP method, the performance can be further improved with only 75% of its parameters.
Generalization Performance.
4.3 Enhancement of State-of-the-Art Method
First, we remove the prompt components of PromptIR. In fact, PromptIR w/o Prompt is identical to Restormer [56], which serves as the baseline for PromptIR. The average improvement of PromptIR over PromptIR w/o Prompt is 0.38 dB on In-Dis, which aligns with the results reported in the original PromptIR paper [41]. Then we directly apply sequential learning (with ‘RJBSNHL’ sequence) to PromptIR. With the same training setting and architecture, sequential learning elevates the PSNR of PromptIR by approximately 0.75 dB on In-Dis and Out-Dis, while obtaining almost the same performance on Uknown. When applying sequential learning and explicit prompt learning (by replacing the prompt components of PromptIR), the performance of PromptIR are improved by over 1 dB on In-Dis and Out-Dis, while being increased by 0.17 dB on Uknown. Note that, PromptIR with our explicit prompt (including classifier) requires only 26.7M parameters, which is 75% of that used in the original PromptIR. These results demonstrate the effectiveness and generality of our learning strategies across various methods.
@@ 4.3 Enhancement of State-of-the-Art Method
Model Params. In-Dis Avg. Out-Dis Avg. Uknown Avg.
SRResNet 1.2M 29.52 24.88 18.02
SwinIR 11.6M 31.05 26.09 18.36
Restormer 26.1M 31.50 25.92 19.30
Uformer 50.9M 30.70 25.62 19.08
PromptIR 35.4M 31.88 26.30 19.38
PromptIR(Restormer)-S+EP (Ours) 26.7M 32.98 27.30 19.55
Table 7: The results of common backbones (SRResNet, SwinIR, Restormer, Uformer) and the recent method PromptIR under our MiO IR formulation. The result of PromptIR coupled with our sequential learning and explicit prompt learning strategies (i.e., “-S+EP”) is also given. The best and second best results are marked in red and blue.
@@ 4.4 Results of More Backbones
As can be seen from Tab. 6, Restormer and Uformer exhibit similar behaviors to SRResNet and SwinIR in the main paper. Sequential and prompt learning can improve the models’ performance on almost all test sets and they could supplement each other. Explicit prompt learning is good at In-Dis and Out-Dis test sets, while adaptive prompt learning is adept at Unknown test sets. However, there are a couple of exceptional cases for Restormer coupled with AP on In-Dis and Out-Dis, where the performance is lower than baseline method Restormer-M. There can be two reasons. First, Restormer contains different structures (e.g., U-like layers and transformer), making it relatively difficult to train. In its original paper [56], a specialized training method is used to train it. Second, Restormer uses channels as tokens to calculate attention, making it difficult to learn tasks that are globally inconsistent (e.g., Dehazing). It always has poor performance on the Dehazing task compared with other methods (see Tab. 8 and Tab. A.1 in the Appendix). Nonetheless, ‘Restormer+AP’ could still perform better than its baseline on the Unknown test set, validating the effectiveness of our proposed strategies.
@@ 4.6 Benchmark MiO IR Models
In Tab. 7, we summarize the results of common backbones (SRResNet, SwinIR, Restormer, Uformer) and the recent method PromptIR under our MiO IR formulation. These methods represent the performance of different “mixed learning (-M)” models on MiO IR. We also show the result of PromptIR coupled with our sequential learning and explicit prompt learning strategies (i.e., ‘-S+EP’) as reference. From the comparison, we could have some findings. For example, though SRResNet lags behind other networks by 1$\sim$2 dB, it only has 1/10$\sim$1/15 the parameters of them. In many resource-constrained situations, the CNN backbone networks can be critical for deployment. Our strategies can steadily improve these networks of varying scales. In addition, it can be found that the network with larger number of parameters (e.g., Uformer) does not necessarily achieve better performance of MiO IR. Therefore, how to design better backbones for MiO IR is also a promising direction to be further explored.
@@ 5 Conclusion
Table 8: Detailed PSNR results on In-Dis test sets. ‘-M’ and ‘-S’ mean mixed and sequential learning, and ‘-EP’ and ‘-AP’ mean explicit and adaptive prompt learning, respectively. ‘-S+EP’ means using sequential learning and explicit prompt learning together, and so on.
@@
SwinIR-S 19.94 25.58 28.16 29.84 34.74 24.33 24.79 26.77 +1.89
SwinIR-M+EP 20.23 25.24 28.01 29.81 34.51 23.30 24.77 26.55 +1.67
SwinIR-S+EP 20.07 25.71 28.30 30.15 35.75 24.10 24.72 26.97 +2.09
SwinIR-M+AP 20.45 24.88 26.99 29.14 31.75 22.18 23.69 25.58 +0.70
SwinIR-S+AP 20.02 25.73 28.20 29.94 35.00 24.17 24.98 26.86 +1.98
@@
Table 9: Detailed PSNR results on Out-Dis test sets. ‘-M’ and ‘-S’ mean mixed and sequential learning, and ‘-EP’ and ‘-AP’ mean explicit and adaptive prompt learning, respectively. ‘-S+EP’ means using sequential learning and explicit prompt learning together, and so on.
@@
Table 10: Detailed PSNR results on Unknown test sets. ‘-M’ and ‘-S’ mean mixed and sequential learning, and ‘-EP’ and ‘-AP’ mean explicit and adaptive prompt learning, respectively. ‘-S+EP’ means using sequential learning and explicit prompt learning together, and so on.
@@
In this Appendix, we present the detailed training / testing degradation settings of MiO IR in Sec. A. Then we provide the results of backbone networks trained on single IR tasks in Sec. B. In addition, we also show more visual results on Out-Dis and Unknown test sets in Sec. C.
A Degradation Settings of MiO IR

### 17 Mar 2024 v2

@@ 1 Introduction
There are some “all-in-one”  [^1] We believe the term of “multiple-in-one” is more precise and appropriate than “all-in-one”. methods  [25, 40, 43] that attempt to handle multiple IR tasks with a single model. However, their setups and explorations have some limitations. Firstly, many of them only take 3 to 5 IR tasks into consideration. Such a small number of tasks cannot reflect the training conflict between different tasks. Secondly, many of the “all-in-one” methods simply adopt the datasets from individual IR tasks in training and testing. However, the ground-truth (GT) images of single task datasets have uneven image quality. These datasets may be suitable for the training of single tasks, but they will mislead the training and evaluation of “all-in-one” IR networks (Some examples are provided in Appendix). These limitations affect the exploration of the research problems along this line. Therefore, despite the progress made by the above methods, by far it remains a challenging issue on how to train a single model to effectively handle multiple IR tasks.
To solve the above mentioned problem, we propose the formulation of Multiple-in-One (MiO) IR, which aims to tackle multiple IR tasks using a single model. There are two pivotal challenges in MiO IR: diverse objective optimization and task adaptation. We then develop two effective and complementary strategies – sequential learning and prompt learning – attempting to address these two challenges, respectively. Specifically, we consider 7 popular IR tasks, including super-resolution, deblurring, denoising, deJPEG, deraining, dehazing and low-light enhancement, and train a single model to handle them all. Compared with the setting in previous works [25, 40, 43], the proposed MiO setup employs high-quality GT images to generate the training and testing data, avoiding the risk of low-quality supervision signals. This formulation also allows us to explore the unique challenges of MiO IR.
@@
For the challenge of task adaptation, we propose the prompt learning  [^2] Prompt learning is similar to conditional learning in this work. Please refer to the related work section for more discussions. strategy. An appropriate prompt can help the network better understand the task at hand and adjust the direction of reconstruction. We provide two methods of prompt learning. One uses additional input as prompt to obtain the task type explicitly (like that in  [40, 33]), while the other adaptively extracts dynamic visual prompt from the input image. These two methods represent the two extreme cases of prompt learning, and are favourable to different application scenarios. As shown in Fig. 1, together with sequential learning, the explicit prompt learning improves the average PSNR by 0.84/1.21 dB for SRResNet/SwinIR, respectively, while the adaptive prompt learning achieves an improvement of 0.24/0.95 dB for SRResNet/SwinIR, respectively. It is worth mentioning that, unlike previous approaches [25, 61], our adaptive prompt learning strategy does not require any specially designed supervision, and its higher generalization ability can be witnessed by an average PSNR improvement of 1.07/0.62 dB for SRResNet/SwinIR across five out-of-distribution test sets. Besides, our strategies can also enhance the state-of-the-art method PromptIR [43] by 1.1 dB with only 75% of its parameters.
@@ 2 Related Work
Prompt learning is originally known from the research on how to introduce additional texts (i.e., prompts) as inputs to pre-trained large language models so that the desired outputs can be obtained [44, 3]. With further research, it becomes common to use different forms of prompts in model training or fine-tuning [2, 52]. In IR field, ProRes [40] and PromptGIP [33] employ additional input images or image pairs as prompts to tell the model what the IR task is. These methods can be viewed as explicit prompt learning. However, in real-world IR applications, sometimes it is difficult to explicitly assign an exact task type for the given image. So it is anticipated to extract information adaptively from the input image as prompt [25, 50]. PromptIR [43] utilizes a classifier-based architecture to extract degradation details from images. However, it requires additional context regarding image degradation, which positions PromptIR close to explicit prompt learning. In this work, both explicit prompt learning and adaptive prompt learning are investigated for the proposed MiO IR formulation.
@@ 3.3 Prompt Learning
For explicit prompt learning, we bind a fixed prompt with each IR task, and then train a classifier to select the prompt for the corresponding IR task during training and testing. In addition to using the classifier, we can also manually specify the task type to select corresponding prompt. Since it is not diffucult to training a high-accuracy classifier for IR tasks, both classifier and manually selection can be viewed as informing the network the explicit task types directly. Considering that the difficulty of explicit prompt learning is not very high, it is supposed to have good performance on in-distribution data.
For real-world low-quality images, sometimes even human subjects cannot explicitly tell the IR task type, and hence explicit prompt learning may fail in such cases. The adaptive prompt learning can be a remedy, where we use the input image as visual prompt to extract task type information adaptively. As illustrated in Fig. 2(c), it has the same architecture as explicit prompt learning except that the input image is used as prompt, instead of any additional input. Because we do not impose any additional constraints on the extractor of prompt, adaptive prompt learning models are much more difficult to train than explicit prompt learning. However, once learned, it has better out-of-distribution generalization capability because the network is able to decide what features to extract by itself.
@@ 4 Experiments and Analysis
In this section, we firstly present the implementation details of MiO IR, including training / testing data and training settings, in Sec. 4.1. Then we show the effectiveness of our proposed learning strategies on in/out of distribution test sets in Sec. 4.2, and use our proposed strategies to enhance the existing state-of-the-art method in Sec. 4.3. After that, we apply our strategies to more backbone networks in Sec.4.4, and interpret the effectiveness of our strategies from the perspective of degradation representation in Sec. 4.5. In addition, we summarize the results of common backbone networks on MiO IR in Sec. 4.6 and Sec. 4.7 for easy access and comparison with future methods. Finally, in Sec. 4.8 we show that by adjusting the prompts in explicit prompt learning, the restoration style can be adjusted.
@@ 4.2 Effectiveness of Learning Strategies
First, we evaluate the effectiveness of sequential learning on the In-Dis test group. As depicted in Tab. 3, compared with mixed learning (‘-M’), basic sequential learning (‘-S’) can enhance the average PSNR of SRResNet/SwinIR by 0.29/0.85 dB across the 7 tasks. It is worth mentioning that, ranging from 0.18 dB (SRResNet on denoising task) to 2.31 dB (SwinIR on deraining task), the performances on all the 7 tasks using both the two backbones are improved, even the least trained task (i.e., low-light enhancement). Compared to mixed learning, sequential learning changes the training data distribution, allowing some tasks to be trained more. However, our experiments validate that the improvement stems mainly from the better optimization rather than merely altering the training data distribution across different tasks because the performances of all tasks are improved.
@@
PromptIR-S+AP (Ours) 26.3M 31.19 -0.69 25.69 -0.61 20.14 +0.76
Table 5: PSNR results of state-of-the-art method PromptIR [43] with and without our learning strategy on In-Dis, Out-Dis and Uknown. We bold the suitable test sets for different strategies. All models are trained under our MiO IR formulation. Our sequential learning could directly enhance PromptIR. By replacing the original prompt method of PromptIR with our EP method, the performance can be further improved with only 75% of its parameters.
@@
Generalization Performance.
4.3 Enhancement of State-of-the-Art Method
First, we remove the prompt components of PromptIR. In fact, PromptIR w/o Prompt is identical to Restormer [58], which serves as the baseline for PromptIR. The average improvement of PromptIR over PromptIR w/o Prompt is 0.38 dB on In-Dis, which aligns with the results reported in the original PromptIR paper [43]. Then we directly apply sequential learning (with ‘RJBSNHL’ sequence) to PromptIR. Because sequential learning is to tackle the optimization problem, it can work in conjunction with PromptIR, which aims at the adaption problem. With the same training setting and architecture, sequential learning elevates the PSNR of PromptIR by approximately 0.75 dB on In-Dis and Out-Dis, while obtaining almost the same performance on Uknown.
When applying sequential learning and explicit prompt learning (by replacing the prompt components of PromptIR), the performance of PromptIR are improved by over 1 dB on In-Dis and Out-Dis, while being increased by 0.17 dB on Uknown. When coupled with sequential learning and adaptive prompt learning, the performance of PromptIR is improved by 0.76 dB on Uknown but drops on In-Dis and Out-Dis. This is because adaptive learning is more suitable for out-of training distribution scenarios. Note that, PromptIR with our explicit prompt (including classifier)/adaptive prompt requires only 26.7M/26.3M parameters, which is 75% of that used in the original PromptIR. These results demonstrate the effectiveness and generality of our learning strategies across various methods.
@@ 4.4 Results of More Backbones
As can be seen from Tab. 6, Restormer and Uformer exhibit similar behaviors to SRResNet and SwinIR in the main paper. Sequential and prompt learning can improve the models’ performance on almost all test sets and they could supplement each other. Explicit prompt learning is good at In-Dis and Out-Dis test sets, while adaptive prompt learning is adept at Unknown test sets. However, there are a couple of exceptional cases for Restormer coupled with AP on In-Dis and Out-Dis, where the performance is lower than baseline method Restormer-M. There can be two reasons. First, Restormer contains different structures (e.g., U-like layers and transformer), making it relatively difficult to train. In its original paper [58], a specialized training method is used to train it. Second, Restormer uses channels as tokens to calculate attention, making it difficult to learn tasks that are globally inconsistent (e.g., Dehazing). It always has poor performance on the Dehazing task compared with other methods (see Tab. 8 and Tab. B.1 in the Appendix). Nonetheless, ‘Restormer+AP’ could still perform better than its baseline on the Unknown test set, validating the effectiveness of our proposed strategies.
Model Params. FLOPs In-Dis Avg. Out-Dis Avg. Uknown Avg.
SRResNet 1.2M 39.98G 29.52 24.88 18.02
SwinIR 11.6M 405.63G 31.05 26.09 18.36
Restormer 26.1M 77.44G 31.50 25.92 19.30
Uformer 50.9M 43.35G 30.70 25.62 19.08
PromptIR 35.4M 86.36G 31.88 26.30 19.38
Restormer-S+EP (Ours) 26.7M 77.90G 32.98 27.30 19.55
Restormer-S+AP (Ours) 26.4M 78.09G 31.19 25.69 20.14
Table 7: The results of common backbones (SRResNet, SwinIR, Restormer, Uformer) and the recent method PromptIR under our MiO IR formulation. The result of PromptIR coupled with our sequential learning and explicit prompt learning strategies (i.e., “-S+EP”) is also given. FLOPs are calculated in $128\times 128$ images. The best and second best results are marked in red and blue. Note that the backbone of PromptIR is Restormer, and thus Restormer-X+XX is equivalent to PromptIR-X+XX.
@@ 4.5 Degradation Representation Analysis
Figure 5: Image restoration style adjustment of Restormer-S+EP by interpolating the prompts of Low-light enhancement and Deraining.
@@ 4.6 Benchmark MiO IR Models
In Tab. 7, we summarize the results of common backbones (SRResNet, SwinIR, Restormer, Uformer) and the recent method PromptIR under our MiO IR formulation. These methods represent the performance of different “mixed learning (-M)” models on MiO IR. We also show the result of PromptIR coupled with our sequential learning and explicit prompt learning strategies (i.e., ‘-S+EP’) as reference. From the comparison, we could have some findings. For example, though SRResNet lags behind other networks by 1$\sim$2 dB, it only has 1/10$\sim$1/15 the parameters of them. In many resource-constrained situations, the CNN backbone networks can be critical for deployment. Our strategies can steadily improve these networks of varying scales. In addition, it can be found that the network with larger number of parameters (e.g., Uformer) or larger number of FLOPs (e.g., SwinIR) does not necessarily achieve better performance of MiO IR. Therefore, how to design better backbones for MiO IR is also a promising direction to be further explored.
@@ 4.7 Detailed Results on Each IR Task
4.8 Restoration Style Adjustment by Adjusting Prompts
By adjusting the prompts of explicit prompt learning, we can adjust the style of restoration outputs. It is even possible to interpolate different prompts to obtain different restoration styles. As shown in Fig 5, with the interpolation of the prompts between low-light enhancement and rain removal, we can adjust the weight of the network for dark light recovery and rain removal. The input image in the first row only contains rain artifacts, which are gradually removed as the weight of the rain removal prompt is increased. The input image in the second row is a low-light image. With the increasing weight of the low-light enhancement prompt, the image is gradually enhanced. The input image in the last row contains both of the two degradations. We can also adjust the effect of rain removal and low-light enhancement by adjusting the prompt weights.
@@ 5 Conclusion
Table 8: Detailed PSNR results on In-Dis test sets. ‘-M’ and ‘-S’ mean mixed and sequential learning, and ‘-EP’ and ‘-AP’ mean explicit and adaptive prompt learning, respectively. ‘-S+EP’ means using sequential learning and explicit prompt learning together, and so on. Note that the backbone of PromptIR is Restormer, and thus Restormer-X+XX is equivalent to PromptIR-X+XX.
@@
SwinIR-S 19.94 25.58 28.16 29.84 34.74 24.33 24.79 26.77 +0.68
SwinIR-M+EP 20.23 25.24 28.01 29.81 34.51 23.30 24.77 26.55 +0.46
SwinIR-S+EP 20.07 25.71 28.30 30.15 35.75 24.10 24.72 26.97 +0.88
SwinIR-M+AP 20.45 24.88 26.99 29.14 31.75 22.18 23.69 25.58 -0.51
SwinIR-S+AP 20.02 25.73 28.20 29.94 35.00 24.17 24.98 26.86 +0.77
@@
Table 9: Detailed PSNR results on Out-Dis test sets. ‘-M’ and ‘-S’ mean mixed and sequential learning, and ‘-EP’ and ‘-AP’ mean explicit and adaptive prompt learning, respectively. ‘-S+EP’ means using sequential learning and explicit prompt learning together, and so on. Note that the backbone of PromptIR is Restormer, and thus Restormer-X+XX is equivalent to PromptIR-X+XX.
@@
Table 10: Detailed PSNR results on Unknown test sets. ‘-M’ and ‘-S’ mean mixed and sequential learning, and ‘-EP’ and ‘-AP’ mean explicit and adaptive prompt learning, respectively. ‘-S+EP’ means using sequential learning and explicit prompt learning together, and so on. Note that the backbone of PromptIR is Restormer, and thus Restormer-X+XX is equivalent to PromptIR-X+XX.
@@
In this Appendix, we first explain in Sec. A why simply adopting the existing single task IR datasets is inappropriate for MiO IR model training. Then we present the detailed training / testing degradation settings of MiO IR in Sec. B and the results of backbone networks trained on single IR tasks in Sec. C. Finally, we show more visual results on Out-Dis and Unknown test sets in Sec. D.
A The Problem of Single Task IR Datasets for MiO IR Model Training
As described in the main paper, there are some “all-in-one” IR methods, which adopt the datasets from single-task IR methods in model training. We argue that this may not be appropriate for the MiO IR model training, because the ground-truth (GT) images in those single-task IR datasets may have degraded quality, and thus the results may be biased for the MiO IR research.
Let’s use the tasks of DeJPEG and Deraining as an example to illustrate the problem. As shown in Fig B.1, the GT images from Rain1200 [59] contain obvious JPEG artifacts. By using this dataset to train a single task Deraining model, the trained model will remove rain but retain JPEG artifacts. However, the MiO model aims to remove the rain and JPEG artifacts simultaneously. As a result, the MiO model will yield better image quality but lower PSNR value on the Rain1200 dataset, because the GT images used to calculate the PSNR metric have JPEG artifacts. Such a problem exists in a few single-task IR datasets that are used in previous “all-in-one” works, such as Rain1200 [59], Rain1400 [16], RESIDE [24], etc. These datasets are sufficient for training and evaluating single IR tasks, but are not appropriate for multiple-in-one IR tasks.
B Degradation Settings of MiO IR